# Having Second Thoughts? Let’s hear it

### v1

@@ 4 Discussion
Our STcert is inspired by top-down attention of the brain, and the current method is one of the simplest possible ways to mimic top-down attention. In this study, we use the same model for both original and second thought predictions to focus on testing the utility of STCert. We expect, however, a secondary model can be optimized further to process the extracted ROIs more effectively, and this custom secondary model can enhance STCert. As the extracted ROIs can be provided by the DL models trained on the original datasets, training the secondary model will not require new datasets. Further, it may be possible to leverage more common senses to certify DL models’ predictions. By combining large language models and foundation segmentation models, we may obtain more thorough information regarding predicted objects and use it to evaluate original predictions. For instance, large language models could automatically recall that automobiles have wheels, and then we could use text prompt-based segmentation models to determine their existence in the scene. We plan to test these approaches in future studies.

### v2

@@ 4 Discussion
Our STcert is inspired by top-down attention of the brain, and the current method is one of the simplest possible ways to mimic top-down attention. In this study, we use the same model for both original and second thought predictions to focus on testing the utility of STCert. We expect, however, a secondary model can be optimized further to process the extracted ROIs more effectively, and this custom secondary model can enhance STCert. As the extracted ROIs can be provided by the DL models trained on the original datasets, training the secondary model will not require new datasets. Further, it may be possible to leverage more contexts and knowledge to certify DL models’ predictions. For instance, by combining large language models and foundation segmentation models, we may obtain more thorough information regarding predicted objects and use it to evaluate original predictions. For instance, large language models could automatically recall that automobiles have wheels, and then we could use text prompt-based segmentation models to determine their existence in the scene. We plan to test these approaches in future studies.
4.1 Limitations
Finally, we list a few major limitations of STCert. First, we underline that STCert is an error detection algorithm and does not have any means to correct incorrect decisions. Consequently, we need additional processes/algorithms to build robust DL models (i.e., DNNs), which can reliably make correct decisions. Similarly, STCert cannot work as a common defense algorithm that enables DL models to make correct predictions/decisions even with adversarial perturbations. Second, STCert is designed for computer vision models. Extending it to other domains may be challenging because STCert requires the identification of proper contexts, and contexts are not well defined in many domains (e.g., natural language processing). Third, STCert heavily relies on segmentation models, which require large computing resources and data for training. Similarly, its inference also demands more computing power than original DL models alone. This issue may partly be mitigated by utilizing a more lightweight segmentation model trained for narrow tasks, when necessary. Fourth, the second prediction was made using a cropped image significantly different from common (ImageNet) training images, which may reduce the accuracy of the second prediction. This issue may be mitigated by training the second classifiers on ROIs explicitly, which we plan to test in the future.
4.2 Broader Impacts
We propose an error detection algorithm for DL vision models. As this algorithm aims to reduce DL models’ mistakes, it could lead to building more reliable DL models. We are not aware of any potential negative impact.
Appendix A Appendix and Supplementary Material
mixed_10 | mixed_13 | living_9 | |||||||
top 1 |
top 2 |
top 3 |
top 4 |
top 5 |
big_12 | geirhos_16 | ||||||||
top 1 |
top 2 |
top 3 |
top 4 |
top 5 |
Model | R18 | R50 | V19 |
top 1 | butcher shop | butcher shop | go-kart |
top 2 | howler monkey | canoe | butcher shop |
top 3 | canoe | confectionery store | canoe |
top 4 | cockroach | stingray | confectionery store |
top 5 | dining table | mountain bike | dining table |
top 6 | confectionery store | dining table | southern black widow |
top 7 | go-kart | cockroach | dung beetle |
top 8 | fireboat | kit fox | cockroach |
top 9 | apiary | football helmet | apiary |
top 10 | dung beetle | fireboat | howler monkey |
Model | DEN | RNEXT | VIT |
top 1 | butcher shop | butcher shop | yellow garden spider |
top 2 | canoe | canoe | canoe |
top 3 | apiary | howler monkey | howler monkey |
top 4 | dung beetle | confectionery store | butcher shop |
top 5 | confectionery store | fireboat | cockroach |
top 6 | European garden spider | football helmet | confectionery store |
top 7 | howler monkey | cockroach | dung beetle |
top 8 | mountain bike | southern black widow | tarantula |
top 9 | cockroach | bookstore | harvestman |
top 10 | fireboat | dining table | boathouse |